OpenStudy (istim):

A yogurt company estimates that the revenue from selling x containers of yogurt is 4.5x. Its cost, C, in dollars, for producing this # of containers of yogurt is C(s)=0.0001x^2+2x+3200. What is the selling price of a container of yogurt?

OpenStudy (istim):

This one I have not done the work for. Relates to my previous question.

OpenStudy (istim):

Seems to relate to revenue formula, which is 4.5x. answer is $4.50. Explanation is the important factor here.

OpenStudy (anonymous):

Total revenue is given by the formula, R = P * Q, where R is total revenue, P is price of good and Q is quantity of good sold. So comparing the equation, R = P * Q with the given R = 4.5 x, where x is the quantity of the yogurt sold. x is equivalent to Q, thus 4.5 is the selling price of the yogurt. Hope this explanation is good enough.

OpenStudy (anonymous):

@tiaph is correct. They threw the cost formula in just to throw you off, it is an easier question than it looks.